OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B is a square pyramid with a base side length of 24 inches and a height of 16 inches. How many times bigger is the volume of pyramid B than pyramid A?

OpenStudy (nincompoop):

show your attempt at solving this on your own

OpenStudy (anonymous):

Area of the base * Height * 1/3 V(A) = 12*8*1/3 V(A)=31.97 V(B)=24*16*1/3 V(B)=127.9 Not sure what to do after this part.

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

The formula for pyramid's volume is (LxWxH) /3 in this case length and width is the same because of the square base

OpenStudy (anonymous):

fix the volumes of pyramids A and B. Then divide the bigger volume by the smaller volume. Lemme know if you got the answer : D

OpenStudy (anonymous):

V(A) = 12^2*8*1/3 V(A)=383.96 V(B) = 24^2*16*1/3 V(B) = 3,071.97 3,071.97/383.96= 8

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

Yes, great! :)

OpenStudy (anonymous):

V(A) is 8 times smaller than v(b) Though I suggest you round answers to 384 and 3072